OpenStudy (sleepyjess):

I know some posts have already been made but......... Congratulations @AriPotta for reaching 99SS

OpenStudy (aripotta):

\(\Huge :D\)

OpenStudy (aripotta):

thank you very much :3

OpenStudy (confluxepic):

Congratulations. @AriPotta @~

OpenStudy (sleepyjess):

Oh, yes I forgot Calvin @~

OpenStudy (confluxepic):

I liked the suspense. @sleepyjess

OpenStudy (sleepyjess):

hehe

TheSmartOne (thesmartone):

this is like the 5th post.

OpenStudy (aripotta):

lol i would have enjoyed the suspense too, but she tagged me in it so i knew it was about me :P

TheSmartOne (thesmartone):

But I made the 1st one. WOhoo!

OpenStudy (sleepyjess):

i know lol

OpenStudy (sleepyjess):

I was supposed to make the 1st post :(

TheSmartOne (thesmartone):

And Congrats!!

OpenStudy (sleepyjess):

AND YOU TOOK THAT AWAY FROM ME!!!!!!

TheSmartOne (thesmartone):

I hope she doesn't go back to 98 like what happened when she reqached 90SS :/

OpenStudy (aripotta):

ugh that probably will happen unfortunately

TheSmartOne (thesmartone):

lol if you see your graph... LOL

OpenStudy (sleepyjess):

THEN I CAN MAKE THE FIRST POST!

TheSmartOne (thesmartone):

lol lets get our subscores up!

OpenStudy (sleepyjess):

i just need to take away some of those medals i gave her and......

OpenStudy (sleepyjess):

lol

OpenStudy (confluxepic):

Calm down. @sleepyjess @!

OpenStudy (aripotta):

i know :| it looks like a roller coaster lol

OpenStudy (confluxepic):

Calvin's family. @~ @! @* @, @;

OpenStudy (sleepyjess):

Calvin has a large family

OpenStudy (confluxepic):

Fish families are cute.

OpenStudy (sleepyjess):

Whenever I see ~ now I think of a space because of LaTeX lol

OpenStudy (aripotta):

only 5 fish in the school. it's not that big. him, his dad, his mum, his sister, and his grandma who has been "just visiting" for a few months now

TheSmartOne (thesmartone):

I need like 76 medals in OS feedback to get the 2nd title... :/

OpenStudy (sleepyjess):

I need 111 more for math

OpenStudy (aripotta):

i need a lot lol

OpenStudy (aripotta):

oh wait. for the third i mean, not second

OpenStudy (sleepyjess):

oh lol

OpenStudy (aripotta):

i need 1874 :O

OpenStudy (sleepyjess):

OMG I HAVEN'T BROUGHT LATEX INTO THIS POST YET

OpenStudy (sleepyjess):

\(\huge\color{red}{\bigstar}\color{orange}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{green}{\bigstar}\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}\\\huge\color{magenta}{Congratulations!}\\\huge\color{red}{\bigstar}\color{orange}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{green}{\bigstar}\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}\)

OpenStudy (sleepyjess):

There we go, better :)

OpenStudy (aripotta):

lol!

OpenStudy (sleepyjess):

It need more LaTeX

OpenStudy (sleepyjess):

\(\Huge\frak{\color{red}H\color{green}a\color{red}p\color{green}p\color{red}y~\color{green}H\color{red}o\color{green}l\color{red}i\color{green}d\color{red}a\color{green}y\color{red}s\color{green}!}\)

OpenStudy (aripotta):

\[\sqrt{(x _{2}-x _{1})^2+(y _{2}-y _{1})^2}\]my latex contribution

OpenStudy (sleepyjess):

lol

TheSmartOne (thesmartone):

^

OpenStudy (confluxepic):

That's the distance formula.

OpenStudy (aripotta):

yes, indeed it is

OpenStudy (aripotta):

more latex:\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (confluxepic):

Slope formula.

OpenStudy (confluxepic):

\[\Huge Slope=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (sleepyjess):

\(l*w*h\)

OpenStudy (confluxepic):

Volume formula.

OpenStudy (sleepyjess):

yep

OpenStudy (sleepyjess):

lol

OpenStudy (sleepyjess):

you have to keep the last one in your profile forever

OpenStudy (sleepyjess):

LOL

OpenStudy (confluxepic):

\[\Huge V-E+F=2\]

OpenStudy (confluxepic):

Euler's Equation.

TheSmartOne (thesmartone):

lol

TheSmartOne (thesmartone):

that is what happened to my post :P

Nnesha (nnesha):

now 35m ago smartone post on hour ago still open both ::(:( :( :(

OpenStudy (sleepyjess):

:D

Nnesha (nnesha):

\(\huge\color{purple}{congratulation!!!}\)you lucky

Nnesha (nnesha):

and i was kidding @AriPotta

OpenStudy (confluxepic):

\[\Huge\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}}\\ \end{array}\]

OpenStudy (confluxepic):

Quadratic Formula.

Nnesha (nnesha):

i guess this is feedback section not math :/ O_O o_O O.o O>O O.o o.o

TheSmartOne (thesmartone):

lol

OpenStudy (confluxepic):

We are math lovers.

TheSmartOne (thesmartone):

@Nnesha I have closed mine now.. to post a new one :P

Nnesha (nnesha):

new one about what ???

TheSmartOne (thesmartone):

something.. just wait :P

OpenStudy (sleepyjess):

wut?

OpenStudy (confluxepic):

^^

OpenStudy (aripotta):

i think she's speaking Klingon

OpenStudy (aripotta):

oh she deleted it :x

TheSmartOne (thesmartone):

lol

OpenStudy (yanasidlinskiy):

Oh, @AriPotta Congrats!!!!

OpenStudy (aripotta):

thank you! \(\color{magenta}{\heartsuit}\)